OpenStudy (narissa):

the graph of a line has a slope of 3 and passes through the point (-4, -5). what is the equation of the line in point slope form?

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (narissa):

@imqwerty

OpenStudy (narissa):

y-4=3(x-5)?

pooja195 (pooja195):

2 negative signs next to each other make a postitive Also, (x1,y1)

pooja195 (pooja195):

-4 is x1 not y1 -5 is y1 not x1

OpenStudy (narissa):

y+5=3(x+4)?

pooja195 (pooja195):

Perfect :)

OpenStudy (narissa):

thank you:))

pooja195 (pooja195):

yw ^_^